OpenStudy (anonymous):

Which three ordered pairs complete the table to give solutions of the equation y = x – 5? A. (–3, –2), (0, –5), and (3, 2) B. (–3, –2), (0, –5), and (3, –2) C. (–3, –8), (0, –5), and (3, 2) D. (–3, –8), (0, –5), and (3, –2)

myininaya (myininaya):

see which 3 order pairs work by plugging the ordered pairs into the equation y=x-5

myininaya (myininaya):

Example. Is (-3,-2) an ordered pair? No because -3-5 is -8 not -2 So this actually implies (-3,-8) is an ordered pair. Check the others.

OpenStudy (anonymous):

is it C.

myininaya (myininaya):

Is (0,-5) an ordered pair on y=x-5? Is (3,2) an ordered on y=x-5?